# Distributional Vision Language Alignment by Cauchy-Schwarz Divergence

### v2

@@ 4 Experiments
We evaluate our method on two tasks to illustrate its vision-language alignment ability: text-to-image (T2I) generation in Section 4.1 and image-text retrieval in Section 4.2. Additionally, we provide the image-text classification results in the Appendix G.
@@ 4.1 Text to image generation
Datasets. Following a previous T2I approach [15], we train our method on four datasets: MSCOCO [16], CC3M [28], CC12M [29], and LAION-HighResolution-5M [30]. MSCOCO contains 80K images paired with multiple captions. CC3M and CC12M include about 2.5M and 10M image-text pairs, respectively. LAION-HighResolution comprises 175M high-resolution pairs, from which we select 5M for training. We evaluate the aligned model on the MSCOCO 30K validation set.
Experimental setup. We build our method based on unCLIP-style approaches (e.g., DALL-E-2 [3], Karlo [31], Kandinsky [4]). These methods typically train a diffusion prior module on large-scale datasets (hundreds of millions data) to map text representations into the image representation space. The diffused text representations are fed into a separate decoder for image generation.
Differently, CS-Aligner trains an adapter to align text representations to image feature space on small-scale datasets, e.g., MSCOCO (0.08M), CC3M (3M), and CC12M (12M), and LAION-HighRes subset (5M). After alignment, we directly process the aligned text features using the pretrained decoder of the large-scale methods (e.g., Karlo and Kandinsky) to generate images, without additional prior modules or multiple diffusion steps. We evaluate generation quality with the FID score [32], which measures how closely generated images match the real image distribution. This metric is particularly well-suited for evaluating modality alignment, as it directly reflects the distribution distance. Additional details can be found in Appendix F.
Table 1: Comparisons with T2I methods. Our method outperforms both large-scale diffusion-based methods (Karlo, Kandinsky) and the recent small-scale method Eclipse.
@@
Baselines. Our baselines consists of both large-scale methods Karlo [31], Kandinsky [4] and recent small-scale alignment method Eclipse [15]. Eclipse streamlines the prior module in Karlo and Kandinsky by employing an $L_{2}$ loss for T2I. For fairness, we use the same Transformer adapter as Eclipse and only align the “CLS” tokens, highlighting the advantages of our distributional alignment.
Comparisons. We compare our method with both the large-scale diffusion-based methods and the small-scale alignment methods. The results are provided in Table 1. By aligning text representations to image representations on the small MSCOCO data, our method achieves superior T2I generation than the large-scale methods, Karlo and Kandinsky, without any diffusion steps. CS-Aligner also outperforms Eclipse by an obvious margin using either Karlo or Kandinsky decoders. The results demonstrate the effective vision-language alignment capability of our method. Moreover, we compare CS-Aligner with Eclipse across different training datasets. As shown in Table 2, our method performs better across diverse training data (CC3M, CC12M, and LAION-HighRes-5M), underscoring the importance of the modality distribution information for robust alignment.
@@
Qualitative Visualization. To further test our method, Fig. 4(a) shows qualitative visualizations of generated images using Karlo decoder. Our aligned text representations result in more realistic images with stronger semantic consistency with the input sentence, highlighting the effectiveness of CS-Aligner in enhancing alignment. More visualizations are provided in Appendix E.1.
@@
CS-Aligner with different adaptation approaches. To demonstrate the robustness of our method across different models, we perform alignments for T2I using both adapter and LoRA. Specifically, we apply LoRA with a low-rank dimension of 8 to every transformer layer in the CLIP text encoder. As shown in Table 3, based on either Karlo or Kandinsky, CS-Aligner with LoRA introduces fewer parameters, while still achieving comparable results compared with the adapter-based one, demonstrating the effectiveness and adaptability of CS-Aligner across different models.
CS-Aligner with multiple captions. It is common in real-world datasets for a single image to correspond to multiple captions (e.g., 5 captions per image in MSCOCO). Due to their pairwise alignment nature, previous methods such as InfoNCE and $\ell_{2}$-based approaches [6, 15] struggle to simultaneously leverage multiple captions. In contrast, by incorporating CS divergence, our CS-Aligner enables training for alignment with single image and multiple captions through the divergence term. To demonstrate the benefits of multiple captions for CS-Aligner, we conducted experiments on the MSCOCO dataset by estimating the CS divergence term $\widehat{D}_{\text{CS}}$ in Eq. (6) using both single and multiple captions. As shown in Fig. 5(a), CS-Aligner effectively leverages the information provided by multiple captions, leading to improved vision-language alignment.
@@
CS-Aligner with additional unpaired data. Collecting and accurately annotating paired vision-language data is both challenging and costly. Enhancing alignment with additional unpaired data offers a more flexible and scalable solution for real-world applications. However, similar to the case of multiple captions, previous methods [6, 15] struggle to fully utilize unpaired data due to their reliance on pairwise alignment, whereas CS-Aligner naturally incorporates the unpaired data information by CS divergence. To demonstrate this capability, we conduct experiments on the MSCOCO dataset using the Kandinsky decoder with (1) 80K paired training samples, (2) 40K paired training samples, and (3) 40K paired training samples supplemented with 80K unpaired samples, where the unpaired samples are used to estimate the CS divergence. As shown in Fig. 5(b), the result with 40K paired training data is lower than 80K. However, introducing additional unpaired data obviously improves the performance, even surpassing the model trained with 80K paired samples. This demonstrates CS-Aligner’s ability to effectively leverage the distributional information of modalities for alignment.
(a) Align with multiple captions.
@@
CS-Aligner with token alignment. Beyond the unpaired data, CS-Aligner also enables token-level alignment by treating the tokens of each sample as a distribution. We evaluated the token-level extension of CS-Aligner using the Kandinsky decoder on the MSCOCO dataset. As shown in Fig. 4(b), incorporating token alignment further improves performance. Moreover, qualitative results indicate that token alignment enhances fine-grained details in generated images, suggesting an improved ability to capture fine-grained relationships between modalities. Additional visualizations are provided in Fig. 7 in Appendix E.2.
@@ 4.2 Image-Text Retrieval
Experimental Setup. Effective multimodal alignment also benefits cross-model retrieval. To demonstrate the alignment ability of our method on retrieval tasks, we align LLMs [25] text representations with CLIP vision representations on both image-to-text and text-to-image retrieval. We use the Flickr 1K test set [33] for short-text retrieval, while Urban1K [34] and DOCCI [35] are employed for long-text retrieval. We compare CS-Aligner against pure InfoNCE-based methods, such as Long-CLIP [34] and LLM2CLIP [1], as the baselines. To ensure a fair comparison, we adopt the setup from LLM2CLIP, aligning CLIP ViT-L/14 image representations with Llama 3 (8B) text representations. Both the vision and text representations are aligned by adapters trained on the CC3M dataset.
@@ 5 Related work
Vision-language alignment and applications. CLIP [6] serves as a foundational model for vision-language alignment in multimodal tasks. Several works have enhanced CLIP through techniques such as momentum distillation [36] and noisy text supervision [37]. Despite its success, CLIP suffers from a persistent modality gap between text and image representations. Prior studies [10, 11, 12] attribute this gap to factors such as cone effects [11] and suboptimal latent space structures [12]. To address this, various strategies have been proposed, including projection adapters [10, 13, 1], geodesic multimodal mixup [14], and parameter-efficient fine-tuning [38]. Recent works also improve CLIP by large language models (LLMs) [39, 2, 1] for downstream tasks such as image-text retrieval.
In addition to image-text retrieval, text-to-image (T2I) generation is another application that reflects the vision-language alignment capability. T2I has advanced significantly over the past decades, driven by both diffusion-based [40, 41, 42, 43] and GAN-based models [44, 45]. Among diffusion-based methods, the unCLIP framework [40, 3] employs a two-stage architecture with a CLIP-guided diffusion prior and a decoder (e.g., DALL-E-2 [3] or Karlo [31]). Its prior module $g_{\phi}$ maps text representations $\mathbf{y}$ to image ones $\mathbf{x}$ by a diffusion model. Recently, Eclipse [15] employs an $\ell_{2}$ loss to simplify the prior loss by eliminating diffusion time and intruding a noise $\epsilon$ term: $\mathcal{L}_{\text{prior}}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\mathbf{x}-g_{\phi}(\epsilon,\mathbf{y})\right\|_{2}^{2}\right].$ However, these methods still rely on pairwise loss (e.g., $\ell_{2}$). In contrast, our approach introduces distributional alignment for a more holistic modality alignment.
Cauchy-Schwarz (CS) divergence. CS divergence [23, 18] is derived from the Cauchy-Schwarz inequality for square-integrable functions. It serves as a symmetric distribution distance metric with notable properties, such as the ability to measure conditional distributions [46] and the closed-form expression for mixtures of Gaussians [47]. CS divergence has been successfully applied across various domains, including deep clustering [48], disentangled representation learning [49], and deep regression [50]. Moreover, due to its advantage of estimating discrepancy between conditional distributions, it has demonstrated success in the domain adaption area [51] and time series clustering [46]. However, the utility of CS divergence in foundation models remains unclear and unexplored.
@@ 6 Conclusion
Limitation & future work. Though our method has numerous applications, it is currently evaluated only on unCLIP-type models for generation. In future, we will explore its integration with stable-diffusion-based models. Broader modalities (e.g., audio and video) and diverse tasks (e.g., image-to-text generation) could further strengthen alignment and broaden the applicability of CS-Aligner.
Appendix A Broader Impacts
@@
Appendix B Details of the toy examples
Example B.1.
Consider two Gaussian distributions, $p(\mathbf{x})\sim\mathcal{N}(\mu_{\mathbf{x}},\sigma_{\mathbf{x}}^{2})$ and $p(\mathbf{y})\sim\mathcal{N}(\mu_{\mathbf{y}},\sigma_{\mathbf{y}}^{2})$, with a joint distribution $p(\mathbf{x},\mathbf{y})\sim\mathcal{N}\left(\begin{pmatrix}\mu_{\mathbf{x}}\\ \mu_{\mathbf{y}}\end{pmatrix},\begin{pmatrix}\sigma_{\mathbf{x}}^{2}&\rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}\\ \rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}&\sigma_{\mathbf{y}}^{2}\end{pmatrix}\right)$. Here, $\mu_{\mathbf{x}}$ and $\mu_{\mathbf{y}}$ are the means of $\mathbf{x}$ and $\mathbf{y}$, $\sigma_{\mathbf{x}}^{2}$ and $\sigma_{\mathbf{y}}^{2}$ are their variances, and $\rho$ is the correlation coefficient and controls their linear dependency. When $\rho=0.99$, the two modalities are highly dependent, with high mutual information ($I=1.959$; see Fig. 2a and 2b). When $\rho=0$, the modalities are independent, resulting in zero mutual information (Fig. 2c). Interestingly, two distributions with the same mutual information value can either exhibit minimal statistical distance and nearly identical shapes, including similar locations, widths, and higher-order moments, as shown in Fig. 2b, or have completely different shapes with distinct means (00 for $p(\mathbf{x})$ and $2$ for $p(\mathbf{y})$) and variances ($4$ for $p(\mathbf{x})$ and $1$ for $p(\mathbf{y})$), as illustrated in Fig. 2a. Quantitatively, the former case shows a minimal KL divergence of 00, while the latter exhibits a KL divergence of nearly $6.81$.
@@ Appendix B Details of the toy examples
[…]
@@
Appendix C Derivations
In this section, we provide a derivation of alignment and uniformity terms of InfoNCE. More concrete analysis can be found in [22].
Let $(\mathbf{x},\mathbf{y})$ be positive (image–text) pairs drawn from $p_{\text{pair}}$, and let $\{(\mathbf{x}^{\prime}_{i},\mathbf{y}^{\prime}_{i})\}_{i=1}^{M}$ be $M$ negative samples (unpaired samples) drawn i.i.d. from the marginal $p_{\text{data}}$. The one-sided InfoNCE (CLIP) loss with temperature $\tau>0$ is

### v3

@@ 4 Experiments
We evaluate our method on two tasks to illustrate its vision-language alignment ability: text-to-image (T2I) generation in Section 4.1 and image-text retrieval in Section 4.2. Note that we focus on the vision-language alignment and use the generation task as a proxy to measure it. Additionally, we provide the image-text classification and the image captioning results in Appendix H. We also present the computation complexity and stability analysis in Appendix E, and additional ablation studies in Appendix H.1.
@@ 4.1 Text to image generation
Datasets. Following a previous T2I approach (Patel et al., 2024), we train our method on four datasets: MSCOCO (Lin et al., 2014), CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), and LAION-HighResolution-5M (Schuhmann et al., 2022). MSCOCO contains 80K images paired with multiple captions. CC3M and CC12M include about 2.5M and 10M image-text pairs, respectively. LAION-HighResolution comprises 175M high-resolution pairs, from which we select 5M for training. We evaluate the aligned model on the MSCOCO 30K validation set.
Experimental setup. We build our method based on unCLIP-style approaches (e.g., DALL-E-2 (Ramesh et al., 2022), Karlo (Donghoon et al., 2022), Kandinsky (Razzhigaev et al., 2023)). These methods train a diffusion prior module on large-scale datasets (hundreds of millions of samples) to map text into the image representation space, and use a decoder to generate images.
Table 1: Comparisons with T2I methods. Our method outperforms large-scale diffusion-based methods and the recent small-scale (alignment) methods (Eclipse and IB (Almudévar et al., 2025)).
@@
Ours + SD-unclip decoder $\text{0.08}_{\text{(COCO)}}$ 10.88
Differently, CS-Aligner trains an adapter to align text representations to image feature space on small-scale datasets, e.g., MSCOCO (0.08M), CC3M (3M), and CC12M (12M), and LAION-HighRes subset (5M). After alignment, we directly process the aligned text features using the pretrained decoder of the large-scale methods (e.g., Karlo and Kandinsky) to generate images, without additional prior modules or multiple diffusion steps. We evaluate generation quality with the FID score (Heusel et al., 2017), which measures how closely generated images match the real image distribution. This metric is particularly well-suited for evaluating modality alignment, as it directly reflects the distribution distance. Additional details can be found in Appendix G.
Baselines. Our baselines consists of both large-scale methods Karlo, Kandinsky, Wurstchen (Pernias et al., 2023), Stable Diffusion (Rombach et al., 2022) (SD v2.1 and SD-unClip), and the recent small-scale alignment method Eclipse. We also compare with the most recent multimodal alignment method (Almudévar et al., 2025) (denoted as IB) on the generation task. For fairness, we use the same Transformer adapter as Eclipse (also for (Almudévar et al., 2025)) and only align the “CLS” tokens, highlighting the advantages of our distributional alignment.
@@
Comparisons. We compare our method with both the large-scale diffusion-based methods and the small-scale alignment methods. The results are provided in Table 1. By aligning text representations to image representations on the small MSCOCO data, our method achieves superior T2I generation than the large-scale methods, Karlo, Kandinsky, and Stable Diffusion without any diffusion steps. CS-Aligner also outperforms Eclipse and IB by an obvious margin using either Karlo or Kandinsky decoders. The results demonstrate the effective vision-language alignment capability of our method. Moreover, we compare CS-Aligner with Eclipse across different training datasets. As shown in Table 2, our method performs better across diverse training data (CC3M, CC12M, and LAION-HighRes-5M), underscoring the importance of the modality distribution information for robust alignment.
Qualitative Visualization. To further test our method, Fig. 4(a) shows qualitative visualizations of generated images using Karlo decoder. Our aligned text representations result in more realistic images with stronger semantic consistency with the input sentence, highlighting the effectiveness of CS-Aligner in enhancing alignment. More visualizations are provided in Appendix F.1.
@@
CS-Aligner with different adaptation approaches. To demonstrate the robustness of our method across different models, we perform alignments for T2I using both adapter and LoRA. Specifically, we apply LoRA with a low-rank dimension of 8 to every transformer layer in the CLIP text encoder. As shown in Table 3, based on different decoders, CS-Aligner with LoRA introduces fewer parameters, while still achieving comparable results compared with the adapter-based one, showing the effectiveness and adaptability of CS-Aligner across different models.
CS-Aligner with multiple captions. It is common in real-world datasets for a single image to correspond to multiple captions (e.g., 5 captions per image in MSCOCO). Due to their pairwise alignment nature, previous methods such as InfoNCE and $\ell_{2}$-based approaches (Radford et al., 2021; Patel et al., 2024) struggle to simultaneously leverage multiple captions. In contrast, by incorporating CS divergence, our CS-Aligner enables training for alignment with single image and multiple captions through the divergence term. To demonstrate the benefits of multiple captions for CS-Aligner, we conducted experiments on the MSCOCO dataset by estimating the CS divergence term $\widehat{D}_{\text{CS}}$ in Eq. (6) using both single and multiple captions. As shown in Fig. 5(a), CS-Aligner effectively leverages the information provided by multiple captions, leading to improved vision-language alignment.
(a) Align with multi-captions.
@@
CS-Aligner with additional unpaired data. Collecting and accurately annotating paired vision-language data is both challenging and costly. Enhancing alignment with additional unpaired data offers a more flexible and scalable solution for real-world applications. However, similar to the case of multiple captions, previous methods (Radford et al., 2021; Patel et al., 2024) struggle to fully utilize unpaired data due to their reliance on pairwise alignment, whereas CS-Aligner naturally incorporates the unpaired data information by CS divergence. To demonstrate this capability, we conduct experiments on the MSCOCO dataset using the Kandinsky decoder with (1) 80K paired training samples, (2) 40K paired training samples, and (3) 40K paired training samples supplemented with 80K unpaired samples, where the unpaired samples are used to estimate the CS divergence. As shown in Fig. 5(b), the result with 40K paired training data is lower than 80K. However, introducing additional unpaired data obviously improves the performance, even surpassing the model trained with 80K paired samples. This demonstrates CS-Aligner’s ability to effectively leverage the distributional information of modalities for alignment.
CS-Aligner with token alignment. Beyond the unpaired data, CS-Aligner also enables token-level alignment by treating the tokens of each sample as a distribution. We evaluated the token-level extension of CS-Aligner with the Kandinsky decoder on MSCOCO. As shown in Fig. 4(b), incorporating token alignment further improves performance. Moreover, qualitative results indicate that token alignment enhances fine-grained details in generated images, suggesting an improved ability to capture fine-grained relationships between modalities. Additional visualizations are provided in Fig. 7 in Appendix F.2.
@@ 4.2 Image-Text Retrieval
Experimental Setup. Effective multimodal alignment also benefits cross-modal retrieval.
@@
To demonstrate the alignment ability of our method on retrieval tasks, we align LLMs (Dubey et al., 2024) text representations with CLIP vision representations on both image-to-text and text-to-image retrieval. We use the Flickr 1K test set (Young et al., 2014) for short-text retrieval, while Urban1K (Zhang et al., 2025) and DOCCI (Onoe et al., 2025) are employed for long-text retrieval. We compare CS-Aligner against pure InfoNCE-based methods, such as Long-CLIP (Zhang et al., 2025) and LLM2CLIP (Huang et al., 2024), as the baselines. To ensure a fair comparison, we adopt the setup from LLM2CLIP, aligning CLIP ViT-L/14 image representations with Llama 3 (8B) text embeddings. Both the vision and text representations are aligned by adapters trained on CC3M.
@@ 5 Related work
Vision-language alignment and applications. CLIP (Radford et al., 2021) serves as a foundational model for vision-language alignment in multimodal tasks. Several works have enhanced CLIP through techniques such as momentum distillation (Li et al., 2021) and noisy text supervision (Jia et al., 2021). Despite its success, CLIP suffers from a persistent modality gap between text and image representations. Prior studies (Zhou et al., 2023; Liang et al., 2022; Shi et al., 2023) attribute this gap to factors such as cone effects (Liang et al., 2022) and suboptimal latent space structures (Shi et al., 2023). To address this, various strategies have been proposed, including projection adapters (Zhou et al., 2023; Gao et al., 2024; Huang et al., 2024), geodesic multimodal mixup (Oh et al., 2024), and parameter-efficient fine-tuning (Zanella and Ben Ayed, 2024). Recent works also improve CLIP by large language models (LLMs) (Jang et al., 2024; Koukounas et al., 2024; Huang et al., 2024) for downstream tasks such as image-text retrieval.
In addition to image-text retrieval, text-to-image (T2I) generation is another application that reflects the vision-language alignment capability. T2I has advanced significantly over the past decades, driven by both diffusion-based (Ramesh et al., 2021; Rombach et al., 2022; Saharia et al., 2022; Nichol et al., 2021) and GAN-based models (Zhang et al., 2017; Tao et al., 2023). Among diffusion-based methods, the unCLIP framework (Ramesh et al., 2021; 2022) employs a two-stage architecture with a CLIP-guided diffusion prior and a decoder (e.g., DALL-E-2 (Ramesh et al., 2022) or Karlo (Donghoon et al., 2022)). Its prior module $g_{\phi}$ maps text representations $\mathbf{y}$ to image ones $\mathbf{x}$ by a diffusion model. Recently, Eclipse (Patel et al., 2024) employs an $\ell_{2}$ loss to simplify the prior loss by eliminating diffusion time and introducing a noise $\epsilon$ term: $\mathcal{L}_{\text{prior}}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|\mathbf{x}-g_{\phi}(\epsilon,\mathbf{y})\right\|_{2}^{2}\right].$ However, these methods still rely on pairwise loss (e.g., $\ell_{2}$). In contrast, our approach introduces distributional alignment for a more holistic modality alignment.
@@ 6 Conclusion
Use of Large Language Models (LLMs).
We used LLMs solely for minor language polishing. They were not involved in research ideation, experimental design, or substantive manuscript writing.
Ethics statement
@@
Reproducibility Statement
We provide sufficient details for reproducibility in Sections 3 and G.
Appendix A Details of the toy examples
Example A.1.
Consider two Gaussian distributions, $p(\mathbf{x})\sim\mathcal{N}(\mu_{\mathbf{x}},\sigma_{\mathbf{x}}^{2})$ and $p(\mathbf{y})\sim\mathcal{N}(\mu_{\mathbf{y}},\sigma_{\mathbf{y}}^{2})$, with a joint distribution $p(\mathbf{x},\mathbf{y})\sim\mathcal{N}\left(\begin{pmatrix}\mu_{\mathbf{x}}\\ \mu_{\mathbf{y}}\end{pmatrix},\begin{pmatrix}\sigma_{\mathbf{x}}^{2}&\rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}\\ \rho\sigma_{\mathbf{x}}\sigma_{\mathbf{y}}&\sigma_{\mathbf{y}}^{2}\end{pmatrix}\right)$. Here, $\mu_{\mathbf{x}}$ and $\mu_{\mathbf{y}}$ are the means of $\mathbf{x}$ and $\mathbf{y}$, $\sigma_{\mathbf{x}}^{2}$ and $\sigma_{\mathbf{y}}^{2}$ are their variances, and $\rho$ is the correlation coefficient and controls their linear dependency. When $\rho=0.99$, the two modalities are highly dependent, with high mutual information ($I=1.959$; see Fig. 2a and 2b). When $\rho=0$, the modalities are independent, resulting in zero mutual information (Fig. 2c). Interestingly, two distributions with the same mutual information value can either exhibit minimal statistical distance and nearly identical shapes, including similar locations, widths, and higher-order moments, as shown in Fig. 2b, or have completely different shapes with distinct means (0 for $p(\mathbf{x})$ and $2$ for $p(\mathbf{y})$) and variances ($4$ for $p(\mathbf{x})$ and $1$ for $p(\mathbf{y})$), as illustrated in Fig. 2a. Quantitatively, the former case shows a minimal KL divergence of 0, while the latter exhibits a KL divergence of nearly $6.81$.
@@ Appendix A Details of the toy examples
$I(\mathbf{x};\mathbf{y})\,=\,\int\!\!\int p(\mathbf{x},\mathbf{y})\,\log\!\Bigl(\tfrac{p(\mathbf{x},\mathbf{y})}{p(\mathbf{x})\,p(\mathbf{y})}\Bigr)\,d\mathbf{x}\,d\mathbf{y}.$ (13)
@@
Appendix B Derivations
In this section, we provide a derivation of alignment and uniformity terms of InfoNCE. More concrete analysis can be found in (Wang and Isola, 2020).
Let $(\mathbf{x},\mathbf{y})$ be positive (image–text) pairs drawn from $p_{\text{pair}}$, and let $\{(\mathbf{x}^{\prime}_{i},\mathbf{y}^{\prime}_{i})\}_{i=1}^{M}$ be $M$ negative samples (unpaired samples) drawn i.i.d. from the marginal $p_{\text{data}}$. The one‐sided InfoNCE (CLIP) loss with temperature $\tau>0$ is